OpenStudy (anonymous):

solve the following quadratic equation by factoring x^4-13x^2+36=0

OpenStudy (anonymous):

(x^2-9)(x^2-4)=0 either x^2-9=0 x^2=9 x=3 ,x=-3 or x^2-4=0 x^2=4 x=2,x=-2

OpenStudy (anonymous):

can you explain please